A solution on the shelf is not labeled for concentration but it is hydrochloric acid. If it’s pH is checked and found to be 0.4, what is its concentration?
Answer #1

Solution:-

Given​​​​​​​ : pH = 0.4

[H​​​​​​+] = 10-pH

[H​​​​​​​+] = 10-(0.4) = 0.398 M

[HCl] = [H​​​​​​​+] = 0.398 M

------------------------------------------